OpenStudy (anonymous):

Use basic identities to simplify the expression. sin^2θ + tan^2θ + cos^2θ

OpenStudy (ipwnbunnies):

These are the two you will need: \[\sin^2 \theta + \cos^2 \theta = 1\] \[1 + \tan^2 \theta = \sec^2 \theta\]

OpenStudy (perl):

hint, use pythagorean identities

OpenStudy (anonymous):

So my answer would be cos4x? @iPwnBunnies

OpenStudy (ipwnbunnies):

No, try again. Are all of your quantities squared? Is your problem this: Simplify \[\sin^2 \theta + \tan^2 \theta + \cos^2 \theta\]

OpenStudy (anonymous):

sorry wrong question i meant tan^2 angle

OpenStudy (ipwnbunnies):

So, it is but I wrote above? Use the Pythagorean identities I gave you to simplify it. It'll look easier if you rearranged it: \[\sin^2 \theta + \cos^2 \theta + \tan^2 \theta\]

OpenStudy (anonymous):

so is it tan2θ? thats what i got @iPwnBunnies

OpenStudy (anonymous):

@sidsiddhartha help :(

OpenStudy (sidsiddhartha):

as iPwnBunnies mentioned above 1+tan^2x=sec^2x